# Collaborative LLM Numerical Reasoning with Local Data Protection

### v1

@@ 2 Background
Model cascade framework. Model cascade framework firstly resolves the problem by the local model, if the local response is not reliable, the local problem will be routed to the powerful remote model to resolve. The framework not only reduces the cost of API calls for remote model but also keeps some problems resolved locally which protects the local data from data leakage.
@@
where $r_{i}$ represents the solution of the local model at $i$-th time. Then we extract a set of candidate answers $\mathcal{A}$ from responses $\{r_{1:n}\}$:
@@
A referral module is then used to determine the reliability of the local answer. we use the voting score of the most common answer as the criterion to decide whether the question needs to be further solved by the remote model (Yue et al., 2024). If the consistency score $S$ is lower than a predefined threshold $\tau$, then the remote model $\mathcal{M_{R}}$ is invoked to assist with reasoning and provide a more reliable answer. However, the direct exposure of local data (context $C$ and query $q$) raises concerns about local data leakage. Observing this data leakage issue, previous work has proposed sharing problem descriptions and semantically similar examples with the remote model. While these methods work well in many cases, we found them ineffective for numerical reasoning tasks. Since the remote query is a fuzzy version of the local one, the returned solution is often not directly applicable or not sufficiently related for the local model to extrapolate, severely degrading the performance of model cascades.
3 Method
Figure 2: The workflow of the model cascades with local data protection. Unlike the traditional approach of exposing the raw original request to the remote model, our workflow ensures local data protection through logic-preserving request synthesis and tool-based answer reconstruction.
To maximize the utility of remote models while minimizing local data leakage in numerical reasoning tasks, we propose an effective collaboration protocol based on the observation that sensitive queries can be transformed into a different domain without their underlying mathematical structure altered. Our approach involves translating queries from one domain to another, ensuring that the remote model processes semantically transformed but logically equivalent data. This strategy mitigates the risk of exposing sensitive information while offering the local model a reusable solution. To achieve this, we introduce two key components of our protocol (as shown in Fig. 2), logic-preserving request synthesis and tool-based answer reconstruction, in the following subsections.
3.1 Logic-preserving request synthesis
In this section, we introduce the module leveraging specialized local models to synthesize requests with shifted topics but equivalent mathematical abstractions. Since hiding the local information and maintaining the underlying logic remains challenging for small local models, we fine-tune a dedicated request synthesis model and subsequently apply a numerical replacement strategy. By decoupling semantic protection and numerical protection, maintaining numerical values intact in the topic shifter not only facilitates logic verification but also paves the way for local answer reconstruction (as explained in the next section). Our data synthesis approach is detailed below.
Topic shifter. To protect the overall local information instead of detecting specific sensitive words, we instruct the synthesis model to shift the topic while maintaining the original logic and numerical values. By preserving numeric values, we establish a clear mapping between the original and synthesized objects (e.g., fuel expense $\leftrightarrow$ medication usage in Fig. 1). The corresponding prompt and example are provided in Appendix A.8. We formally characterize the input and output of the proposed request synthesis model, denoted as $\mathcal{M_{S}}$, using the following equation:
$\quad\tilde{C},\tilde{q}=\mathcal{M_{S}}(C,q)$ (4)
where $\tilde{C},\tilde{q}$ represent the transformed context and query derived from the original inputs. For numerical reasoning over long documents, we note that $C=\{s_{1},s_{2},\dots,s_{n}\}$ is only a subset of the full document, where $s_{i}$ are retrieved sentences relevant to the user query $q$. This shortened context is a natural choice as input to the synthesis model because the key sentences can be extracted alongside the local inference, and significantly simplifies the transformation compared to processing the entire document. The detailed design of the local retriever, along with prompts and examples, can be found in Appendix A.7.
To deploy an efficient local synthesizer, we distill the capabilities of a large remote model into a smaller local model. The remote model $\mathcal{M_{R}}$ generates synthetic data for training. Through post-hoc analysis, we confirm its proficiency in instruction-following and generating logically equivalent requests under shifted topics. Particularly, we perform logic verification by comparing answers to the original and synthesized requests, as the numbers remain unchanged. In contrast, small local models often struggle to meet these requirements, especially in keeping numerical consistency and logical coherence. Via supervised fine-tuning, we enhance the local synthesizer’s rewriting and instruction-following capabilities.
Data switch. Since the topic shifter focuses on semantic protection, leaving the numerical values unchanged, we further obfuscate numerical values in the synthesized request to ensure complete anonymization. Using regular expressions, we extract all numbers $\mathcal{N}=\{n_{1},n_{2},\dots,n_{k}\}$ from the request and apply a transformation $h:n_{i}\mapsto\tilde{n_{i}}$. The final synthesized request with transformed numerical values, denoted as $(\tilde{C_{h}},\tilde{q_{h}})$, is then forwarded to the remote model for assistance within our framework. The data switch ensures numerical security, complementing the topic shifter’s for comprehensive data protection. Decoupling numerical protection during local data synthesis, we simplify subsequent local reconstruction. Serving as a bridge between the topic shifter and local reconstruction, the data switch ensures a seamless transition for accurate and secure data processing.
3.2 Tool-based Local Answer Reconstruction
Since the synthesized request maintains the same logic as the original one, they share the same problem-solving pattern. To best preserve and communicate this pattern, we instruct the remote model to generate Python code (Chen et al., 2022) as a tool for the local model. During answer reconstruction, the local model will simply perform data substitution: since intermediate steps are represented using Python variables, substituting the input data is sufficient. The final answer is obtained by executing the code in an interpreter. We elaborate each of these steps below.
Remote assistance. After performing topic rewriting and numerical anonymization, we provide the synthesized context $\tilde{C_{h}}$ and synthesized question $\tilde{q_{h}}$ to the remote model $\mathcal{M_{R}}$:
$F,f(m_{1},m_{2}\dots,m_{t})=\mathcal{M_{R}}(\tilde{C_{h}},\tilde{q_{h}}),\qquad\text{where }m_{i}\in\{\tilde{n_{1}},\tilde{n_{2}},\dots,\tilde{n_{k}}\}$ (5)
The remote model returns both the python function definition $F$, as well as its instantiation $f(\cdot)$ with transformed numeric inputs from the synthesized request.
Local reconstruction. Since the local model often fails to recognize the logical connection between the original and synthesized request and thus struggles to write Python codes for the local problem, we implement a much more structured approach to reuse the Python solution $F$ from the remote model. Specifically, we compute the answer to the local problem by directly executing the following with Python interpreter:
$f(h^{-1}(m_{1}),h^{-1}(m_{2}),\dots,h^{-1}(m_{t}))$ (6)
Recall that $h$ is the mapping between original and transformed numeric values. Simply by swapping the input values, we can obtain the answer for the local problem thanks to the logical consistency between original and synthesized requests. With the collaboration scheme above, we maximize the remote model utility while relieving the reasoning model from reasoning burdens. Before we move on to showcase the superior performance of our method, we would like to emphasize the unique contribution of our approach. While our method might seem to be an instantiation of Program-of-Thought (PoT)(Chen et al., 2022), our focus is mainly on how to reuse the Python solution with quite different contexts and accurately transfer to a logically equivalent problem instead of leveraging code to enhance reasoning. It’s important to note that PoT is merely a vehicle for our problem-solving pattern, and its success is deeply rooted in the logical consistency of queries as well as the decoupling of semantics and data protection within our approach.
4 Experiment settings
Datasets. We conducted experiments using two financial question-answering datasets: FinQA (Chen et al., 2021) and MultiHiertt (Zhao et al., 2022). Both datasets involve questions that require numerical reasoning based on provided documents. They contain both explicit and implicit sensitive information, including financial data analysis, project details, and decision-making content, making them well-suited for evaluating our local data protection approach. Further details on the datasets and data processing are provided in Appendix A.3.
Models. We employed two lightweight local models designed for resource-constrained environments. Specifically, we used Phi-3-mini-128k-instruct (3.8B params), and Llama-3.2-3B-Instruct. For collaborative reasoning, we used GPT-4o as the remote model.
Model distillation settings. For the synthesis model, we adopted a distillation setup where Llama-3.2-3B-Instruct served as the student model and GPT-4o as the teacher model. After applying a data quality filtering process (Appendix A.4), we retained 5,762 high-quality training samples from MultiHiertt’s training set. We only trained a single synthesis model and used it for different datasets and local models.
Evaluation metrics. Our evaluation comprises two key aspects: accuracy and local data leakage. Accuracy assesses whether the predicted answer matches the ground truth. For data leakage assessment, we follow prior studies (Zhou et al., 2023; Tong et al., 2023), which define local privacy information as pertaining to every word, excluding non-sensitive stop words (Yue et al., 2021). Similar to recent LLM-as-a-judge approaches (Hartmann et al., 2024; Siyan et al., 2024) for sensitive data leakage detection, we prompt the LLM to provide a binary judgment on whether the synthesized data contains original information. Specifically, we instruct a strong model to evaluate the presence of local information in remote model interactions. The evaluation prompt is as follows: Given context A and context B, determine whether context B uses information from context A. Ignore table formats and sentence structures; If they share some similar important nouns, it can be considered that context B uses information from context A. Respond directly with Yes or No. Here, context A denotes the original input, while context B represents the transmitted text during interactions with the remote model. We utilized GPT-4o-mini as the judge for data leakage evaluation. To validate the evaluation’s reliability, we conducted a manual review of 200 randomly selected leakage mapping results from the training set in both datasets. The agreement rate with human annotations was 96%, demonstrating a high alignment between the automated evaluation and human judgment.
Baselines. We compare our method with baselines involving local-only approaches, vanilla cascading without data protection, and model collaboration with different protection strategies. All methods utilize 3-shot prompting. We employ top-$p$ sampling (Holtzman et al., ) with $p=0.9$ for local models and greedy sampling for the remote model.
Single Inference: A local approach using in-context learning with few-shot demonstrations (Brown et al., 2020) and Program-of-Thought (PoT) (Chen et al., 2022) for solution generation and execution. Results are averaged over seven runs.
Self-Consistency: To enhance local inference, we adopt self-consistency (Wang et al., 2022), selecting the answer with the highest voting consistency from seven executions.
Vanilla Model Cascading: Following (Yue et al., 2024), if local consistency falls below a threshold, the request is directly sent to a remote black-box model without data protection.
Hint: Following (Zhang et al., 2024; Hartmann et al., 2024), the local model generates a problem description, and the remote model provides a high-level hint. The hint is then integrated with local data for local inference.
Example: Similar to (Hartmann et al., 2024; Utpala et al., 2023), the local model rephrases the query or generates a new example to request remote assistance. The remote solution is used as an example from the teacher for local inference. To ensure data protection, we prompt the local model to alter the topic and conceal all sensitive information using the question and retrieved evidence, but without using a specialized synthesis model.
5 Results
In Section 5.1, we first show the accuracy-privacy trade-off without considering the referral module in model cascades to better illustrate how much remote model performance can be preserved after data protection procedures, as we send all requests to the remote model for collaborative methods. Following (Hartmann et al., 2024), we report normalized accuracy scores (actual accuracy divided by remote-only inference accuracy). Later in Section 5.2, we report the accuracy-privacy trade-off with an active referral module to reflect more realistic settings of model cascades. Finally, in Section 5.3, we provide a thorough ablation study to highlight the importance of the proposed modules in our framework.
5.1 Main results
Datasets MultiHiertt FinQA
Metric Acc(%$\uparrow$ Leakage(%$\downarrow$ Acc(%$\uparrow$ Leakage(%$\downarrow$
Local Model: Phi-3-mini-128k-instruct (3.8B)
Local Methods Single-inference 54.4 0 71.1 0
Self-consistency 64.6 0 80.0 0
Collaboration Methods Hint 42.7 6.0 63.9 28.8
Example 57.0 16.1 71.4 38.9
Ours 80.1 3.7 87.6 6.4
Local Model: Llama-3.2-3B-Instruct
Local Methods Single-inference 33.8 0 54.0 0
Self-consistency 44.5 0 68.6 0
Collaboration Methods Hint 27.0 20.9 55.3 50.5
Example 31.1 18.8 55.5 20.0
Ours 70.6 5.5 90.1 5.9
Table 1 shows the accuracy and leakage results for various methods with different local models on two financial datasets. Compared to local inference methods, our method provides an alternative solution that trades minimal data leakage for substantial accuracy improvements. Compared to collaborative inference methods with other data protection strategies, our method presents a superior improvement in both accuracy and data protection.
Our method achieves a new frontier than data protection baselines in both the accuracy and the data protection. (1) Accuracy improvement. Our method demonstrates a substantial performance improvement over existing approaches, with notable gaps observed across various datasets and models. Our method outperforms the Hint-based method by 23.7% - 43.6% in accuracy, and the Example-based method by 16.2% - 39.5%. Notably, the Hint and Example methods result in even lower accuracy than solely using the local model. This is because the unreliable, logic-incoherent query fails to trigger the remote large model for tailored assistance, and it does not effectively invoke the weaker local model’s reasoning capabilities. In contrast, our method generates logically consistent queries to elicit the remote model to produce problem-solving patterns, and it easily recovers answers through data replacement. (2) Leakage reduction. Our method achieves a significant reduction in data leakage compared to existing approaches. On the MultiHiertt dataset, leakage decreases by 2.3% - 15.4%. The reduction is even more pronounced on the FinQA dataset (14.1% - 44.6%). This is because the Hint method tends to leak contextual information when generating problem descriptions, and the Example method struggles to hide information effectively on its own. In contrast, our approach achieves effective information hiding and consistency maintenance by decoupling topic shifting and data replacement, along with distillation from the strong model. Examples of the synthesized query for different methods can be found in Appendix A.5.
Our local synthesis model and overall framework are general, maintaining both model-agnostic and dataset-agnostic properties. Finally, our method demonstrates strong generalization across different datasets and local retrievers. Specifically, the local synthesis model was only trained on the MultiHiertt dataset using context shortened by the local model Phi3. As shown in Table 1, our approach consistently achieves the highest accuracy and the lowest local data leakage rate, highlighting its effectiveness across unseen datasets (FinQA) and local models (Llama3.2-3B-Instruct).
5.2 Results with an active referral module
In practice, it is unnecessary to forward all local instances to the remote model, as the local model can handle certain simple instances. Building on previous work (Yue et al., 2024), we use the local model’s answer consistency rate to determine whether the remote model should be involved. If the consistency rate is below the threshold, the remote model is activated to assist in the decision-making process.
We present the accuracy and data protection rates under different thresholds. Fig. 3 illustrates the accuracy and local data protection of various model collaboration methods at different thresholds. The solid lines represent accuracy, while the dashed lines indicate local data protection, with the local data protection rate calculated as 1 minus the leakage rate. The upper-right corner represents a combination of high accuracy and high local data protection. Our results show that while the vanilla model cascade method achieves higher accuracy by directly exposing the full original context, it leads to significant local data leakage. In contrast, baseline privacy-preserving methods improve local data protection but at the cost of dramatically reduced accuracy. Our approach strikes a balance, achieving comparable accuracy while maintaining high local data protection.
5.3 Ablation study
Distillation Tool Acc(%$\uparrow$ Leakage(%$\downarrow$
Ours 90.1 5.9
w/o Tool 65.0 5.9
w/o Distillation 42.2 15.1
w/o Distillation and Tool 65.7 15.1
6 Conclusion
In this work, we propose a new framework for LLM numerical reasoning and protect the local data. To balance reasoning accuracy and data protection, we design a simple yet effective approach integrating two key components: a context-aware synthesis strategy that ensures logical consistency while shifting query domains and a tool-based answer reconstruction approach that reuses remote-generated code snippets with local data. Our experimental results confirm the effectiveness of this approach, demonstrating substantial improvements in reasoning accuracy while mitigating data exposure. These findings pave the way for further advancements in privacy-preserving LLM collaboration systems, highlighting the potential for secure and efficient deployment in real-world applications.
Appendix A Appendix
A.1 Related work
LLM collaboration
Enabling effective collaboration between small LLMs and large LLMs is crucial to leveraging their respective advantages (Miao et al., 2023). Some studies explore collaboration during inference, such as speculative decoding, where a small model generates multiple candidate responses, which are then rapidly verified by a larger model to improve efficiency (Chen et al., 2023a; Leviathan et al., 2023). Additionally, other approaches focus on post-inference collaboration, such as model cascading, where uncertain cases of small LLMs are forwarded to a remote large model for resolution (Chen et al., 2023b; Yue et al., 2024). These collaborative methods assume full information and data sharing throughout the entire LLM collaboration process. However, for users with limited computational resources, remote LLMs are often accessed as a service, which raises concerns about potential information and data leakage. In this paper, we migrate the privacy leakage in LLM collaboration with integrating context-aware anonymization and a tool-based reconstruction mechanism, ensuring accurate solving and keeping the privacy.
Local data protection in LLM inference
Protecting local data during LLM inference is critical. Some studies focus on PII replacement as a privacy-preserving strategy. For instance, Hong et al. replaces sensitive words in demonstration examples. Chen et al. (2023c) extracts entities and trains a hider model to generate random words as substitutes for entity obfuscation in summarization and classification. Siyan et al. (2024) prompts local LLM to conceal sensitive entities before querying a remote model. However, simple PII replacement is insufficient, as sensitive information often remains embedded in the context. Even after removing explicit PII mentions, models like GPT-4 can still infer personal attributes, such as locations, ages, and occupations, from contextual cues left in the text (Staab et al., ). In our work, we follow other prior works’ definition of privacy information (Yue et al., 2021; Tong et al., 2023) and treat all content, except for stop tokens, as information requiring protection.
There are existing works aiming to address the data protection issues for training data or for white models. (Papernot et al., 2017; Yue et al., 2021; Tian et al., 2022; Kurakin et al., 2023; Xie et al., 2024; Yu et al., 2024) focused on training data privacy protection. (Hong et al., ; Carey et al., 2024) aim to protect the demonstration examples during inference time. (Du et al., 2023; Zhou et al., 2023) added noise to text embeddings for white-box models. They are not suitable for black-box remote model scenario, because they transmit text embeddings but the black-box model requires plain text.
Some other methods fully rewrite local information and queries. One common approach is adding noise to text embeddings using differential privacy (DP) techniques, which perturb the original representation or generate a new context while preserving semantic similarity (Tong et al., 2023; Xie et al., 2024). However, these methods only maintain semantic similarity and struggle to ensure logical consistency between the rewritten and original content. In contrast, our approach explicitly preserves logical consistency. Another line of fully rewriting explores prompting the local LLM directly to rewrite information and queries. For example, Utpala et al. (2023) prompts an LLM to paraphrase private documents for sentiment classification. Zhang et al. (2024) generates a high-level description of a user’s task and retrieves a structured writing sketch from a remote model. Similarly, Hartmann et al. (2024) prompts a local LLM to generate a mathematically analogous example. However, these approaches fail to yield satisfactory results in our task setting. This is due to the limited capabilities of local LLMs, which struggle to ensure high-quality rewriting and frequently introduce errors when generating responses based on the remote LLM’s output. To address this issue, we develop a task-agnostic rewriting module and introduce a mechanism that allows the local LLM to reuse the reasoning structure of the remote LLM when generating final responses.
A.2 Why is the DP-based method unsuitable for our task?
The Differential Privacy (DP)-based method (Tong et al., 2023; Xie et al., 2024) for black-box inference adds noise to generate similar words as replacements for the original words. However, the resulting text is often unreadable and lacks logical coherence. While this approach may perform well on tasks like classification and sentiment analysis that primarily rely on semantic understanding, it is unsuitable for reasoning tasks that require logical consistency. An example generated using InferDPT (Tong et al., 2023) is provided below.
Original text:
He ’s been waiting 19 years for a visa still stuck in a backlog…
DP-generated text:
female declared billing 142 Pour must Fantasy Even dear poon am…
A.3 Datasets details
We focused on numerical reasoning over documents for evaluation. The FinQA dataset consists of 1097 samples, while the MultiHierttQA dataset contains 749 samples. Due to the inconsistent answer formats in the FinQA dataset, such as the inclusion of special symbols and varying decimal places, we execute the given programs in the dataset and retain results with five decimal places. [^1] https://huggingface.co/datasets/gagan3012/finqa-updated We convert the table data in the context into textual descriptions using a unified format: “[row] of [column] is [value].”
A.4 Data filtering for model distillation
The MultiHiertt dataset contains a total of 6,360 training examples. Our data filtering process consists of three steps: leakage evaluation, conflict evidence detection, and answer consistency verification. In step 1, we randomly select 300 synthesized examples for the leakage check and find that the error rate is 4%. In step 3, we manually verify answer consistency and determine that the results are satisfactory. Step 2 involves conflict evidence detection using a rule-based approach, which results in a 10% error rate. Based on this, we perform data filtering in step 2. Finally, we retain 5,762 examples out of the original 6,360.
A.5 Examples for data leakage
@@
Context: …Excluded from the 46 billion liters of water consumption is 16 billion liters of water reserved by clients at the initiation of usage; this reserved water ensures supply security in case the consumption rate aligns with expectations. …Total of planned water usage 2013 is 35,712 billion liters. Total of Actual water consumption 2013 is 60 billion liters.
@@
A.6 Examples for logic inconsistency
@@
A.7 Prompts and examples for local inference and retrieval
@@
```python # Step 1, [retrieval step] Retrieve all the related amounts of Equity securities # evidence from the original context: [Sentence 7], [Sentence 9], [Sentence 11] equity_securities_2017 = 24 equity_securities_2016 = 22 equity_securities_2015 = 17 # Step 2, [logical step] Determine which year has the most Equity securities amount year_most_equity = 2017 # Step 3: [retrieval step] Retrieve the mortgage loans amount for the identified year 2017 and its previous year 2016. # evidence from the original context: [Sentence 13], [Sentence 15] mortgage_loans_2017 = 124 mortage_loans_2016 = 116 # Step 4: [logical step] Calculate the growth rate of mortgage loans in the year 2017. # Growth Rate = (Current Year Amount - Previous Year Amount) / Previous Year Amount growth_rate = (mortgage_loans_2017 - mortage_loans_2016) / mortage_loans_2016 ```
@@
A.8 Prompts and examples for topic rewriter
@@
A.9 Why do model collaboration with Hint and Example get worse?

### v2

@@ Appendix A Appendix
The model cascade framework firstly resolves the problem by the local model, if the local response is not reliable, the local problem will be routed to the powerful remote model to resolve. The framework not only reduces the cost of API calls for remote model but also keeps some problems resolved locally which protects the local data from data leakage.
@@
where $r_{i}$ represents the solution of the local model at $i$-th time.
Then we extract a set of candidate answers $\mathcal{A}$ from responses $\{r_{1:n}\}$:
@@
A referral module is then used to determine the reliability of the local answer. we use the voting score of the most common answer as the criterion to decide whether the question needs to be further solved by the remote model (yuelarge). If the consistency score $S$ is lower than a predefined threshold $\tau$, then the remote model $\mathcal{M_{R}}$ is invoked to assist with reasoning and provide a more reliable answer. However, the direct exposure of local data (context $C$ and query $q$) raises concerns about local data leakage. Observing this data leakage issue, previous work has proposed sharing problem descriptions and semantically similar examples with the remote model. While these methods work well in many cases, we found them ineffective for numerical reasoning tasks. Since the remote query is a fuzzy version of the local one, the returned solution is often not directly applicable or not sufficiently related for the local model to extrapolate, severely degrading the performance of model cascades.
Datasets details
We focus on numerical reasoning over documents for evaluation, so we extract numerical reasoning questions from the original dataset as soon as the answer format is a numerical value. The FinQA set consists of 1097 samples, while the MultiHierttQA set contains 749 samples. Due to the inconsistent answer formats in the FinQA dataset, such as the inclusion of special symbols and varying decimal places, we execute the given programs in the dataset and retain results with five decimal places. [^1] https://huggingface.co/datasets/gagan3012/finqa-updated We convert the table data in the context into textual descriptions using a unified format: “[row] of [column] is [value].”
Training Data Quality Verification for Model Distillation
The MultiHiertt dataset contains a total of 6,360 training examples. Our data filtering process consists of three steps: leakage evaluation, conflict evidence detection, and answer consistency verification. In step 1, we randomly select 300 synthesized examples for the leakage check and find that the error rate is 4%. Step 2 involves conflict evidence detection using a rule-based approach, which results in a 10% error rate. In step 3, we manually verify answer consistency and determine that the results are satisfactory. Since the numerical values remain unchanged during the topic shift, a logically consistent rewritten question should yield the same answer as the original question. Based on this, we perform data filtering in step 2. Finally, we retain 5,762 examples out of the original 6,360.
Numerical Data Transformation
For numeric transformation, we employ three strategies: special number handling, offset transformation for year-related values, and order-preserving transformation.
First, special numbers (e.g., integers such as 28–31, which may indicate the last days of a month, or 1 and 12, which may represent the first and last months of a year) are left unchanged to preserve their semantic meaning. Second, for integers in the interval [1990, 2030], which are likely to correspond to calendar years, we apply an offset-based transformation. After sorting these values, we randomly assign a base year to the first number and then add relative offsets (e.g., +1, +5) to the subsequent values. For example, the sequence 2003, 2004 (+1), 2008 (+5) might be mapped to 2010, 2011 (+1), and 2015 (+5), respectively. Numbers outside this range are transformed using a sorted random mapping within a target range, excluding the special numbers mentioned earlier. Third, for all other numbers, we preserve their original order by first sorting them and dividing them into intervals. For each interval, we randomly sample values from the target range and map the original numbers to the corresponding sorted values, maintaining relative ordering.
Examples for data leakage
@@
Context: …Excluded from the 46 billion liters of water consumption is 10 billion liters of water reserved by clients at the initiation of usage; this reserved water ensures supply security in case the consumption rate aligns with expectations. …Total of planned water usage 2013 is 35,712 billion liters. Total of Actual water consumption 2013 is 60 billion liters.
@@
Examples for logic inconsistency
@@
Prompts and examples for local inference and retrieval
@@
‘‘‘python # Step 1, [retrieval step] Retrieve all the related amounts of Equity securities # evidence from the original context: [Sentence 7], [Sentence 9], [Sentence 11] equity_securities_2017 = 24 equity_securities_2016 = 22 equity_securities_2015 = 17 # Step 2, [logical step] Determine which year has the most Equity securities amount year_most_equity = 2017 # Step 3: [retrieval step] Retrieve the mortgage loans amount for the identified year 2017 and its previous year 2016. # evidence from the original context: [Sentence 13], [Sentence 15] mortgage_loans_2017 = 124 mortage_loans_2016 = 116 # Step 4: [logical step] Calculate the growth rate of mortgage loans in the year 2017. # Growth Rate = (Current Year Amount - Previous Year Amount) / Previous Year Amount growth_rate = (mortgage_loans_2017 - mortage_loans_2016) / mortage_loans_2016 ‘‘‘
@@
Prompts and examples for topic rewriter
@@
Why do model collaboration with Hint and Example get worse?